OpenStudy (anonymous):

Find the surface area of a square pyramid with length 9 ft and slant height 12 ft.

Directrix (directrix):

Lemme check the formula.

Directrix (directrix):

Okay. Area of Square Pyramid = 2 * e * L + e^2 where e is the length of the base edge, L is the slant height, and e^2 is the area of the base of the pyramid. so ..

Directrix (directrix):

Trouble: Is this posted correctly: square pyramid with length 9 ft. I don't know what the length of a pyramid means. Would you check your problem source?

OpenStudy (anonymous):

It's fine Directrix. (:

Directrix (directrix):

I am going to assume that 9 is the base edge length. (The question is poorly worded) Let's see what pops out.

Directrix (directrix):

Total Area = 2*9*12 + 9^2 @dopegyal_ --> 2*9*12 + 9^2 = ?

Directrix (directrix):

Post what you get and we'll compare answers. @dopegyal_

Directrix (directrix):

These are the answers- ?? A. 90 ft2 B. 45 ft2 C. 171 ft2 D. 297 ft2

OpenStudy (anonymous):

297 is what I got

OpenStudy (anonymous):

YES!(:

Directrix (directrix):

Yes, that is what I got. I was just looking at this same question as asked on OpenStudy before. So, let's go with the 297.

OpenStudy (anonymous):

Do you mind helping me with another question Mr. Einstein ? (: @Directrix

OpenStudy (anonymous):

You surely did. (:

Directrix (directrix):

Yes, post the question here in this thread.